# Vector Quantization using Gaussian Variational Autoencoder

### v1

@@ 4.3 Ablation Studies
Gaussian VAE (w/ TDC) O(1)O(1)/$O(1)$ 32.75 0.023 0.906 0.460
MRC (original) $O(2^{DKL(q(Z_{i}|X)||\mathcal{N}(0,1))})$/$O(1)$ 32.09 0.023 0.906 0.425
ORC $O(2^{DKL(q(Z_{i}|X)||\mathcal{N}(0,1))})$/$O(1)$ 32.09 0.023 0.906 0.419
GQ (Ours) $O(D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)))$/$O(1)$ 32.11 0.023 0.907 0.414
The TDC Parameters and Grouping Strategies See Appendix D.
@@ 5 Related works
VQ-VAE (van2017neural) is an autoencoder that compresses images into discrete tokens. Due to the discretization, it is impossible to be train directly using gradient descent. Various techniques have been proposed to address this, such as commitment loss (van2017neural), expectation maximization (EM) (roy2018theory), the straight-through estimator (STE) (bengio2013estimating), and Gumbel-softmax (jang2016categorical; maddison2016concrete; sonderby2017continuous; Shi2024TamingSV). In addition, VQ-VAE is prone to codebook collapse. To mitigate this, various methods have been proposed, such as reducing the code dimension (Yu2021VectorquantizedIM; Sun2024AutoregressiveMB), product quantization (Zheng2022MoVQMQ), residual quantization (Lee2022AutoregressiveIG), dynamic quantization (Huang2023TowardsAI), multi-level quantization (Razavi2019GeneratingDH), feature projection (Zhu2024ScalingTC), rotation codebooks (Fifty2024RestructuringVQ) and etc. (yu2021vector; chiu2022self; takida2022sq; zhang2023regularized; huh2023straightening; gautam2023soft; goswami2024hypervq).
More related to our work, some variants of VQ-VAE with fixed codebook emerge, such as FSQ (mentzer2023finite), LFQ (yu2023language), and BSQ (zhao2024image). However, training tricks such as the straight-through estimator (STE) is still required. Among all previous works, TokenBridge (wang2025bridging) and ReVQ (zhang2025quantizethenrectifyefficientvqvaetraining) are most aligned with our objective. TokenBridge and ReVQ also convert a pre-trained Gaussian VAE into a VQ-VAE. However, TokenBridge and ReVQ do not constraint the divergence of Gaussian VAE, leading to suboptimal performance. Besides, ReVQ requires some training while our approach is training-free.
Reverse Channel Coding See Section. 3.6.
@@ 6 Conclusion & Discussion
In this paper, we propose Gaussian Quant (GQ), an extremely simple yet effective technique that converts a pre-trained Gaussian VAE into a VQ-VAE without any additional training. Theoretically, we show that when the logarithm of the GQ codebook size exceeds the bits-back coding bitrate of the Gaussian VAE, a small quantization error is achieved. In addition, we propose a target divergence constraint (TDC) to implement GQ effectively. Empirically, we demonstrate that GQ outperforms previous discrete VAEs, such as VQGAN, FSQ, LFQ, and BSQ (van2017neural; mentzer2023finite; yu2023language; zhao2024image). Furthermore, our TDC also improves the performance of TokenBridge (wang2025bridging).
We limit our evaluation of GQ to the standard SD3.5 UNet (Esser2024ScalingRF) and the BSQ-ViT architecture. Additionally, we restrict the bpp range to 0.220.221.001.00, which extends the BSQ’s original range of 0.250.250.500.50 bpp. We acknowledge that there are several highly competitive VQ-VAEs that adopt multi-scale or residual architectures (Tian2024VisualAM; Han2024InfinitySB) or study the low bpp regime (bpp $\leq 0.2$) (Yu2024AnII; Sargent2025FlowTT; zhang2025quantizethenrectifyefficientvqvaetraining). However, in this paper, we use standard architectures and a typical bpp range to focus on the core aspects of the quantization method. Additionally, we focus on achieving a good trade-off between bitrate and reconstruction quality, leaving the complex relationship between reconstruction and generation performance to future works (wang2024image; xiong2025gigatok; hansen2025learnings).
Ethics Statement
@@
Reproducibility Statement
For theoretical results, the proofs for all theorems are presented in Appendix A. For the experiments, we use publicly accessible datasets such as ImageNet (Deng2009ImageNetAL). Implementation details and hyper-parameters are provided in Appendix C. Besides, we include the source code for reproducing the experimental results in the supplementary material.
@@ Appendix A Proof of Main Results
Theorem 1. Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i},\sigma_{i}$, and assuming $|\mu_{i}|,\sigma_{i}$ ’s product and sum are bounded by $|\mu_{i}\sigma_{j}|\leq c_{1}$, $|\mu_{i}|+|\sigma_{j}|\leq c_{2}$, then the probability of quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays double exponentially to the amount of nats $t$ that codebook bitrate $\log K$ exceeds the bits-back coding bitrate, i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-e^{t}\underbrace{\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5}}_{\textup{constant}})}.$ (6)
@@
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-\mu_{i}\sigma_{i}-0.5})}.$ (18)
@@
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (19)
@@
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+\mu_{i}\sigma_{i}-0.5})}$
$\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}$ (20)
@@
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (21)
@@
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-|\mu_{\max}\sigma_{\max}|-0.5})}$
$=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-|\mu_{\max}\sigma_{\max}|-0.5})}$
@@
Theorem 2. The probability of quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increase exponentially to the amount of nats $t$ that codebook bitrate $\log K$ lower than the bits-back coding bitrate, i.e.,
$\textup{when }\log K$ $=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))-t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\geq 1-e^{-t}\underbrace{\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}}_{\textup{constant}}.$ (7)
@@
Similar to the proof of Theorem. 1, we have
@@
$=1-K\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+|\mu_{i}\sigma_{i}|-0.5}$
$\geq 1-Ke^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (25)
@@
$=1-K\sqrt{\frac{2}{\pi}}e^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))+0.5(\mu_{i}^{2}+\sigma_{i}^{2})+|\mu_{i}\sigma_{i}|-0.5}$
$=1-Ke^{-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (26)
Taking the value of $K=e^{D_{KL}(q(Z_{i}||\mathcal{N}(0,1)))-t}$ , we have the result
$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-e^{D_{KL}(q(Z_{i}||\mathcal{N}(0,1)))-t-D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$
@@
To better illustrate the significance of these bounds, we provide a practical example. We evaluate the ImageNet validation dataset using a pre-trained Gaussian VAE and compute that $c_{1}=8.12$ and $c_{2}=1.50$. We then visualize the upper bound and lower bound of $\textup{Pr}{|\hat{z}-\mu_{i}|\geq\sigma_{i}}$ in Fig. 3. The results show that when the codebook bitrate exceeds the bits-back coding bitrate by approximately 10 nats, the probability of large quantization error diminishes to zero. Conversely, when the codebook bitrate is smaller than the bits-back coding bitrate, the probability of large quantization error increases.
Figure 3: A visualization of large quantization error lowerbound and upperbound with ImageNet validation dataset.

### v2

@@ 4.2 Main Results
Diffusion
Gaussian VAE w/o TDC - 8.35 202.19
Gaussian VAE w/ TDC - 8.47 205.94
Auto-regressive
VQGAN 16.4% 8.01 151.40
FSQ 94.3% 7.33 224.88
LFQ 24.9% 7.73 142.09
BSQ 99.8% 7.82 221.64
TokenBridge 94.6% 7.82 198.24
GQ (Ours) 100.0% 7.67 230.79
Table 7: The reconstruction result when $\log K\neq R_{i}$. GQ works best when $\log K$ matches $R_{i}$.
[…]
14 14 $2^{14}\times 1024$ 25.31 0.064 0.762 0.491
18 14 $2^{14}\times 1024$ 25.24 0.068 0.774 0.527
14 18 $2^{18}\times 1024$ 27.79 0.059 0.808 0.513
18 18 $2^{18}\times 1024$ 27.86 0.054 0.804 0.424
Table 8: GQ is robust to random codebook.
[…]
42 27.61 0.059 0.807 0.529
43 27.61 0.059 0.807 0.523
44 27.62 0.059 0.807 0.526
Table 9: The necessity of GQ’s two-stage pipeline is demonstrated: converting a pre-trained Gaussian VAE yields better performance than training GQ from scratch, and further fine-tuning provides only marginal improvements.
[…]
VQ-VAE from scratch 1.00 8.50 0.763 0.156 360
Eq. 10 from scratch 29.65 0.044 0.866 0.928
Convert from Eq. 10 32.47 0.023 0.907 0.327
Further finetune Eq. 10 32.45 0.022 0.905 0.264
Table 10: Ablation study on hyperparameters $\alpha$, $\beta$, and codebook dimension $m$. Here, $\alpha^{\mathrm{bits}}=\alpha/\log 2$.
[…]
0.5 1.01 1 25.60 0.088 0.702 1.26
0.5 1.01 4 26.98 0.068 0.793 0.927
0.5 1.01 16 27.61 0.059 0.807 0.529
0.1 1.01 16 27.56 0.058 0.812 0.551
1.0 1.01 16 27.61 0.063 0.811 0.701
0.5 1.1 16 27.63 0.060 0.809 0.534
0.5 1.001 16 27.48 0.058 0.804 0.510
Image Generation Table 6 evaluates GQ for image generation. Compared with VQGAN, FSQ, LFQ, and BSQ, GQ achieves higher codebook usage and codebook entropy. In terms of generation FID and IS, GQ is comparable to FSQ and outperforms the other methods. Additionally, we train a DiT (Peebles and Xie, 2022) with the same architecture and training setup using Gaussian VAEs with and without TDC. The results show that, under limited computational resources, autoregressive generation is more efficient than diffusion in both FID and IS. This demonstrates that converting a Gaussian VAE to a VQ-VAE facilitates autoregressive generation, improving image generation efficiency.
Complexity Compared with a Gaussian VAE, the computational overhead of GQ is negligible (see Appendix D.6).
4.3 Ablation Study
GQ Conversion: Alternatives Stochastic alternatives to GQ conversion exist, including MRC, ORC, and A coding (Havasi et al., 2018a; Theis and Yosri, 2021; Flamich et al., 2022; He et al., 2024). Table 4 compares these methods in terms of reconstruction quality. When applied to a TDC-constrained Gaussian VAE, GQ achieves the best PSNR, SSIM, and rFID. Moreover, for codebook dimension $m=1$, GQ can be implemented via bisection search, making it asymptotically faster (see Appendix D.7).
GQ Conversion: Selection of Codebook Size Theorem 3.1-3.2 show that setting codebook size $\log K=R_{i}$ is theoretically optimal. In Table 7, we show that empirically GQ is most effective when $\log K$ matches $R_{i}$.
GQ Conversion: Robustness to Random Codebook The codebook $c_{1:K}$ is sampled from a Gaussian distribution and fixed once generated. Table 8 shows that GQ’s performance is nearly invariant to different random seeds.
GQ Training: TDC To evaluate the necessity of TDC in Eq. 9, we train a vanilla Gaussian VAE without TDC. As shown in Table 5, the mean $R_{i}^{\mathrm{bits}}$ of the vanilla VAE is close to that of the TDC-constrained VAE (3.99 vs. 4.26 bits), but the range of $R_{i}^{\mathrm{bits}}$ is much wider for the vanilla model (0.26–27.29 vs. 2.93–5.63 bits). While the reconstruction performance of the two Gaussian VAEs is similar (PSNR: 32.73 vs. 32.61 dB, rFID: 0.490 vs. 0.460), GQ applied to the TDC-constrained VAE significantly outperforms GQ applied to the vanilla VAE (PSNR: 31.25 vs. 26.43 dB, rFID: 0.372 vs. 0.978). This demonstrates that TDC is essential for effective GQ conversion.
Alternatives to TDC include the MIRACLE heuristic (Havasi et al., 2018b) and IsoKL (Flamich et al., 2022). MIRACLE is less effective at controlling the range and is outperformed by TDC (PSNR 29.48 vs. 32.11 dB). IsoKL imposes a stricter constraint by requiring that $R_{i}$ is identical across all dimensions. While IsoKL enforces this constraint effectively, its overall performance is worse (PSNR 30.45 vs. 32.11 dB) due to numerical instability and the exclusion of solutions with $\sigma_{i}^{2}>1$. In Appendix B, we propose a numerically stable version of Mean-KL (Lin et al., 2023), which extends IsoKL to support multi-dimensional codebooks ($m>1$). However, it performs poorly for ViT-based models.
Necessity of GQ’s Two-Stage Pipeline One could train a vanilla VQ-VAE (Van Den Oord et al., 2017) using the same codebook as GQ, effectively creating a VQ-VAE with a fixed Gaussian noise codebook. Alternatively, the Gaussian VAE network could be trained directly with the GQ target in Eq. 10 using Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2016). However, as shown in Table 9, both approaches fail to converge reliably. Moreover, fine-tuning GQ after initializing with a pre-trained Gaussian VAE yields only marginal performance improvements.
Hyperparameters To justify our choice of TDC parameters $\alpha=0.5$ and $\beta=1.01$, we perform a grid search with results shown in Table 10. The results indicate that the effect of $\alpha$ and $\beta$ on TDC is minimal as long as $\alpha\leq 0.5$. Table 10 also shows that a larger codebook dimension $m$ improves reconstruction. Accordingly, we set $m$ to the quotient of the total latent dimension and the total number of tokens, which is the maximum feasible value. To achieve optimal reconstruction, we use $\omega=2.0$ for bitrates $\leq 0.50$ bpp and $\omega=0.0$ for 1.00 bpp (see Table 14 for details).
@@ 5 Related works
Vector-Quantized Variational Autoencoder VQ-VAE (Van Den Oord et al., 2017) is an autoencoder that compresses images into discrete tokens. Due to discretization, it cannot be trained directly via gradient descent. Various techniques address this, including commitment loss (Van Den Oord et al., 2017), expectation maximization (EM) (Roy et al., 2018), the straight-through estimator (STE) (Bengio et al., 2013), and Gumbel-Softmax (Jang et al., 2016; Maddison et al., 2016; Sønderby et al., 2017; Shi et al., 2024b). VQ-VAE is also prone to codebook collapse. To mitigate this, methods such as reducing code dimension (Yu et al., 2021b; Sun et al., 2024b), product quantization (Zheng et al., 2022), residual quantization (Lee et al., 2022), dynamic quantization (Huang et al., 2023), multi-level quantization (Razavi et al., 2019), feature projection (Zhu et al., 2024), and rotation codebooks (Fifty et al., 2024) have been proposed (Yu et al., 2021a; Chiu et al., 2022; Takida et al., 2022; Zhang et al., 2023; Huh et al., 2023; Gautam et al., 2023; Goswami et al., 2024).
More closely related to our work, some VQ-VAE variants have fixed codebooks, including FSQ (Mentzer et al., 2023), LFQ (Yu et al., 2023), and BSQ (Zhao et al., 2024). However, training still relies on tricks such as the straight-through estimator (STE). On the other hand, TokenBridge (Wang et al., 2025) and ReVQ (Zhang et al., 2025) convert a pre-trained Gaussian VAE into a VQ-VAE, but they do not constrain the divergence of the Gaussian VAE.
Reverse Channel Coding GQ is closely related to reverse channel coding, which aims to simulate a distribution $q$ using samples from another distribution $p$ (Harsha et al., 2007; Li and El Gamal, 2018; Havasi et al., 2018a; Flamich et al., 2020; Theis and Yosri, 2021; Flamich et al., 2022; He et al., 2024). The key difference between MRC and GQ is that MRC and its variants (Havasi et al., 2018a; Theis and Yosri, 2021; Flamich et al., 2022; He et al., 2024) simulate a distribution via stochastic sampling, whereas a VQ-VAE requires deterministic quantization. Regarding quantization error, GQ outperforms MRC by construction (Eq. 3). Moreover, GQ with a one-dimensional codebook ($m=1$) can be implemented via bisection search, achieving superior asymptotic complexity (see Appendix D.7).
Additionally, TDC is closely related to the MIRACLE heuristic and IsoKL parameterization of Gaussian VAEs (Havasi et al., 2018b; Flamich et al., 2022; Lin et al., 2023). Specifically, MIRACLE also adjusts $\lambda$ during VAE training, but it maintains a single $\lambda$, making it less effective at controlling the minimum and maximum values of $R_{i}$. IsoKL, in contrast, enforces strict control over $R_{i}$ by directly solving for $\sigma$ given $\mu$ using the Lambert $\mathcal{W}$ function (Corless et al., 1996; Brezinski, 1996), but it suffers from numerical instability and yields suboptimal performance.
@@ 6 Conclusion & Discussion
In this paper, we propose Gaussian Quant (GQ), a simple yet effective method that constructs a VQ-VAE by first training a constrained Gaussian VAE then converting it into a VQ-VAE. The conversion process involves generating codebook from pure Gaussian noise and matching the posterior mean. Theoretically, we show that when the logarithm of the GQ codebook size exceeds the bits-back coding bitrate of the Gaussian VAE, the resulting quantization error is small. To train a constrained Gaussian VAE for effective conversion, we introduce the target divergence constraint (TDC). Empirically, GQ outperforms prior discrete VAEs, including VQGAN, FSQ, LFQ, and BSQ (Van Den Oord et al., 2017; Mentzer et al., 2023; Yu et al., 2023; Zhao et al., 2024), and TDC further improves the performance of TokenBridge (Wang et al., 2025).
We acknowledge that several highly competitive VQ-VAEs employ multi-scale or residual architectures (Tian et al., 2024; Han et al., 2024). In this work, however, we use a standard single-scale architecture to focus on the core aspects of the quantization method. Furthermore, recent studies show that strong reconstruction does not necessarily imply strong generation, and achieving good generation often requires feature alignment (Wang et al., 2024; Xiong et al., 2025; Hansen-Estruch et al., 2025). Here, we focus on reconstruction and leave the exploration of the complex relationship between reconstruction and generation performance to future work.
Impact Statement
@@ Appendix A Proof of Main Results
Lemma A.1.
Given $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1)),|\mu_{i}|,|\sigma_{i}|$ are upper-bounded. And therefore $c_{1},c_{2}$ in Theorem 3.1-3.2 are also bounded:
$|\mu_{i}|\leq\sqrt{2R_{i}},$
$|\sigma_{i}|\leq\sqrt{-\mathcal{W}_{-1}(-e^{-(2R_{i}+1)})},$
$c_{1}\leq\sqrt{2R_{i}}\sqrt{-\mathcal{W}_{-1}(-e^{-(2R_{i}+1)})},$
$c_{2}\leq\sqrt{2R_{i}}+\sqrt{-\mathcal{W}_{-1}(-e^{-(2R_{i}+1)})}.$
Proof.
First, we have
$\mu_{i}^{2}+\sigma_{i}^{2}-1-\log\sigma_{i}^{2}=2R_{i}.$ (15)
Notice that $\sigma_{i}^{2}-\log\sigma_{i}^{2}\geq 0$. To maximize $\sigma_{i}$, we simply set $\mu_{i}=0$. Then we have $\sigma_{i}^{2}-\log\sigma_{i}^{2}=2R_{i}+1$. We can solve
$\sigma_{i}^{2}=-\mathcal{W}(e^{-(2R_{i}+1)}),$ (16)
where $\mathcal{W}(.)$ is Lambert $\mathcal{W}$ function. We select the larger solution
$\sigma^{2}=-\mathcal{W}_{-1}(e^{-(2R_{i}+1)}),$ (17)
where $\mathcal{W}_{-1}$ is the lower branch of Lambert $\mathcal{W}$ function. To maximize $\mu_{i}$, we simply set $\sigma_{i}=1$. Then we have
$\mu_{i}^{2}=2R_{i},\mu_{i}=\sqrt{2R_{i}}.$ (18)
Theorem 3.1. Denote the mean and standard deviation of $q(Z_{i}|X=x)$ as $\mu_{i}$ and $\sigma_{i}$, respectively. We assume $|\mu_{i}\sigma_{i}|\leq c_{1}$ and $|\mu_{i}|+|\sigma_{i}|\leq c_{2}$. Given a fixed $R_{i}=D_{KL}(q(Z_{i}|X)||\mathcal{N}(0,1))$, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ decays doubly exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ exceeds the bits-back coding rate. That is,
$\textup{when }\log K=R_{i}+t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\leq\exp{(-e^{t}\sqrt{\frac{2}{\pi}}e^{-c_{1}-0.5})}.$ (19)
@@
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-\mu_{i}\sigma_{i}-0.5})}.$ (25)
@@
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (26)
@@
$=\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}+\mu_{i}\sigma_{i}-0.5})}$
$\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}$ (27)
@@
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}.$ (28)
@@
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}$ $\leq\exp{(-K\cdot\sqrt{\frac{2}{\pi}}e^{-R_{i}-|\mu_{\max}\sigma_{\max}|-0.5})}$
$=\exp{(-e^{t}\cdot\sqrt{\frac{2}{\pi}}e^{-|\mu_{\max}\sigma_{\max}|-0.5})}.$
@@
Theorem 3.2. Following Theorem 3.1, the probability of a quantization error $|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}$ increases exponentially with the number of nats $t$ by which the codebook bitrate $\log K$ falls below the bits-back coding rate. That is,
$\textup{when }\log K=R_{i}-t,$
$\Pr\{|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i}\}\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5c_{2}^{2}-0.5}.$ (30)
@@
Similar to the proof of Theorem. 3.1, we have
@@
$=1-K\sqrt{\frac{2}{\pi}}e^{-R_{i}+|\mu_{i}\sigma_{i}|-0.5}$
$\geq 1-Ke^{-R_{i}}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (33)
@@
$=1-K\sqrt{\frac{2}{\pi}}e^{-R_{i}+0.5(\mu_{i}^{2}+\sigma_{i}^{2})+|\mu_{i}\sigma_{i}|-0.5}$
$=1-Ke^{-R_{i}}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$ (34)
Taking the value of $K=e^{R_{i}-t}$ , we have the result
$\Pr(|\hat{z}_{i}-\mu_{i}|\geq\sigma_{i})$ $\geq 1-e^{R_{i}-t-R_{i}}\sqrt{\frac{2}{\pi}}e^{0.5(\mu_{i}+\sigma_{i})^{2}-0.5}$
$\geq 1-e^{-t}\sqrt{\frac{2}{\pi}}e^{0.5(|\mu_{i}|+|\sigma_{i}|)^{2}-0.5}$